A chef draws cookies randomly from a box containing 6 cookies of the same shape and size. There is 1 chocolate chip cookie, 3 almond, and 2 butter cookies. He draws 1 cookie and then draws another cookie without replacing the first one. Find the probability of picking one almond cookie followed by another almond cookie and show the equation used.
6 total cookies.....1 choc chip, 3 almond, 2 butter
probability of picking 1 almond on first pick : 3/6 or 1/2
without replacement
probability of drawing another almond : 2/5
Probability of drawing 1 almond followed by another (without replacement ) is:
3/6 * 2/5 = 6/30 = 1/5
Probability of picking two almond cookies = 1 / 5
Probability Formula = Favourable outcomes / Total outcomes
Total outcomes, given = 6 cookies
Pr (first almond cookie) = 3 / 6
Pr (second almond cookie), without replacement = 2 / 5
So, Prob (Two almond cookies) = ( 3 / 6 ) x ( 2 / 5 )
6 / 30 = 1 / 5
To learn more about Probability, refer https://brainly.com/question/9222285?referrer=searchResults